# Convergent Power Series for Anharmonic Chain with Periodic Forcing

### v3

@@ Abstract.
We study the case of a pinned anharmonic chain of oscillators, with coordinates $({\bf q}.{\bf p})=\{(q_{x},p_{x}):\,x\in{\mathbb{Z}}_{N}=\{-N,......,N\}\}$, subjected to an external driving force ${\cal F}(\cdot)$ of period $\theta=2\pi/\omega$ acting on the oscillator at $x=0$. The system evolves according to a Hamiltonian dynamics with frictional damping, $\gamma>0$, present at both endpoints $x=-N,N$. The Hamiltonian is given by
${\cal H}_{N}({\bf q}.{\bf p})=\sum_{x\in{\mathbb{Z}}_{N}}\left[\frac{p_{x}^{2}}{2}+\frac{1}{2}(q_{x}-q_{x-1})^{2}+\frac{\omega_{0}^{2}q_{x}^{2}}{2}+\nu\Big{(}V(q_{x})+U(q_{x}-q_{x-1})\Big{)}\right],$ (0.1)
where $V(\cdot)$ and $U(\cdot)$ are $C^{2}$ smooth anharmonic pinning and interaction potentials with bounded second derivatives, $\omega_{0}>0$ and $\nu\in\mathbb{R}$. We prove that if no integer multiplicity of $\omega$ falls in the spectrum of the infinite harmonic system ${\cal I}:=[\omega_{0},\sqrt{\omega_{0}^{2}+4}]$, then there exists a $\nu_{0}>0$ such that for $|\nu|<\nu_{0}$ the system approaches asymptotically in time a unique $\theta$-periodic solution, whose coordinates are given by a convergent power type series in $\nu$. The series coefficients are bounded and smooth functions of $\nu$. The value of $\nu_{0}$ is a lower bound on the radius of convergence, independent of $N$, the friction coefficient $\gamma>0$ and ${\cal F}$. It depends only on the supremum norm of $V^{\prime\prime}(\cdot)$ and $U^{\prime\prime}(\cdot)$ and the distance of the set of integer multiplicities of $\omega$ from ${\cal I}$.
@@ 1. Introduction
The study of one-dimensional coupled oscillator models out of equilibrium plays an important role in statistical mechanics. On the atomistic level it was initiated by the seminal work of Fermi-Pasta-Ulam-Tsingou [17]. Their work was perhpas the first application of digital computers to a non-military scientific goal. The results of simulations, where they found lack of ergodicity in a chain with anharmonic interactions, are still a source of investigation and a challenge to theoretists, [7, 18, 15, 31]. There has also been much work on transport properties of anharmonic chains both for systems in contact with thermal reservoirs and for systems driven by external periodic forces, see [1, 2, 5, 29, 11, 12, 13, 21, 22, 16, 33, 34] and references therein.
In the present paper we study a one-dimensional anharmonic chain with an external periodic forcing applied at a single site and damping forces at its ends.
Computer simulations on such a system, where a quartic anharmonicity is added to a pinned harmonic potential, reveal a rich variety of steady states emerging as the driving parameters of the external force are varied, see [25, 23, 5, 6, 32, 26, 8, 20].
The harmonic version of this model was rigorously studied by us in a recent work [19]. Here, we demonstrate that, under specific conditions on the anharmonicity and the forcing frequency, a rigorous perturbation scheme can be established around the pure harmonic case.
1.1. Microscopic dynamics
@@
We denote ${\mathbb{Z}}_{N}:=\{-N,\ldots,N\}$. The microscopic dynamics of the chain are given by the forced Hamiltonian system with the Hamiltonian given by (0.1), friction on both endpoints of the lattice interval and a driving periodic force at $x=0$. This yields
@@
with the parameters $\gamma>0,\omega_{0}>0,\theta>0$. The Neumann laplacian $\Delta$ and the discrete gradient $\nabla$ are defined as
@@
We prove that when all integer multiplicities of $\omega=\frac{2\pi}{\theta}$ are outside the interval containing the spectrum of the harmonic chain ${\cal I}:=[\omega_{0},\omega_{u}]$, with $\omega_{u}:=\sqrt{\omega_{0}^{2}+4}$, a unique state of the system is achieved, as $t\to+\infty$, for an arbitrary initial condition. In addition, the state is $\theta$-periodic and given by a perturbative expansion in the parameter $\nu$, that converges for $|\nu|<\nu_{0}$ for all fixed $N$, $\gamma>0$, with
@@
Here $\delta_{*}:=\inf\Big{[}|(m\omega)^{2}-w^{2}|:\,m\in{\mathbb{Z}},\,w\in{\cal I}\Big{]}$, and ${\mathfrak{V}}=\|V^{\prime\prime}\|_{\infty}+3\|U^{\prime\prime}\|_{\infty}$.
Finding a periodic stable solution of (1.2) by a convergent perturbative series is, we believe, impossible for nonlinear anharmonicity that grows faster than quadratic, considered in [3, 28, 15], and, as far as we know, not proven for any other non-equilibrium system. Our results do not extend to the case where some integer multiplicity of $\omega$ lies inside spectrum of the harmonic chain or $|\nu|>\nu_{0}$.
@@ 2.4. Global stability of the unique periodic solution in the anharmonic case
Theorems 2.2 and 2.3 hold even in the absence of dissipation at the boundaries ($\gamma=0$). Of course in this case the unique periodic solution has no stability properties, it is just a very special solution of (1.2). Stability can be proven when $\gamma>0$.
@@ 2.5. Work done by periodic forcing
${\cal W}(\nu)=\frac{1}{\theta}\int_{0}^{\theta}{\cal F}(t/\theta)p_{0,{\rm p}}(t;\nu)\mathrm{d}t.$ (2.18)
@@
Suppose that $\gamma>0$ and $\big{(}q_{x,{\rm p}}(\cdot;\nu)\big{)}$ is a $\theta$-periodic solution of (1.2). Furthermore assume that the potential $U$ is such that the equation [^2] Condition 2.20 means that the interaction potential $\frac{1}{2}r^{2}+\nu U(r)$, between the neighboring atoms does not become ”flat” in some intervals, which would prevent the energy transfer in the chain.
@@
${\cal W}(\nu)=\frac{\gamma}{\theta}\Big{(}\int_{0}^{\theta}p_{-N,{\rm p}}^{2}(s)\mathrm{d}s+\int_{0}^{\theta}p_{N,{\rm p}}^{2}(s)\mathrm{d}s\Big{)}>0.$ (2.21)
@@ 2.6. Spatial decay of the periodic solution as $N\to\infty$
There exists positive constants $C_{1},A,\rho$ depending only on $\omega,\gamma$ and independent of $N$ such that for $|\nu|<\nu_{0}\wedge C_{1}$
@@ 2.7. The case of odd harmonic modes: loss of uniqueness
We stress here the fact that Theorem 2.7 does not claim either the uniqueness, or stability of the solution given by the series (2.15), in the case $\nu_{0}\leqslant|\nu|<\overline{\nu}_{0}$. In this interval the periodic solution exists (2.15) but it may not be stable or unique. Then, see Section 5.1, the system (1.2) may admit more than one $\theta$-periodic solution. Obviously, in that case, the stability in the sense claimed in Theorem 2.4 cannot hold.
@@ 3. The case $N=0$: a single anharmonic oscillator with damping
$\delta_{*}=\left\{\begin{array}[]{ll}\min\{\phi(m_{*}),\phi(m_{*}+1)\},&\mbox{where}\\ &\\ m_{*}:=\left\lfloor\frac{(\omega_{0}^{2}-2\gamma^{2})^{1/2}}{\omega}\right\rfloor&\end{array}\right.$ (3.5)
@@ 4.1. Time harmonics of a periodic solution
00 0 $=\big{[}(\omega m)^{2}-\omega_{0}^{2}+\Delta-i\gamma\omega m(\delta_{x,-N}+\delta_{x,N})\big{]}\widetilde{q}_{x}(m;\nu)+\widehat{\rm F}_{m}\delta_{x,0}$ (4.4)
@@ 4.2. The case of a harmonic chain
[…]
@@
00 0 $=\left[\Delta+(m\omega)^{2}-\omega_{0}^{2}-i\gamma m\omega(\delta_{-N}(x)+\delta_{N}(x))\right]\widetilde{q}_{x}^{(L)}(m)-\;\widetilde{v}_{x,L-1}(m;\nu),\;\quad x\in{\mathbb{Z}}_{N},$ (4.13)
@@ 4.4. Regularity of the periodic solution
00 0 $=\Big{(}\Delta+(m\omega)^{2}-\omega_{0}^{2}-i\gamma m\omega(\delta_{-N}+\delta_{N})\Big{)}\Big{(}\delta\widetilde{Tf}_{x}(m)\Big{)}+$ (4.33)
@@
By the Banach contraction principle there exists a unique solution of the equation ${Tf}=f$ in $L^{2}\big{(}[0,\theta];{\mathbb{R}}^{2N+1}\big{)}$. Then $f(t)$ given by (4.32) is the unique $\theta$-periodic solution of (1.2). Hence the conclusion of the theorem follows. ∎
@@
see (4.14). To see (5.2) (and (5.1)) assume for simplicity that $U\equiv 0$. The argument in the general case is similar. Since $V^{\prime}(\cdot)$ is odd and real valued, its Fourier transform $\widehat{V^{\prime}}(\xi)$ is also an odd function. To further simplify the argument we suppose that its support is compact. The general case can be treated by an approximation argument. Thanks to (4.9) and the fact that $\widehat{F}_{2m}=0$ for all $m$, we conclude that (5.1) holds for $L=0$. Suppose now that (5.2) is true for some non-negative integer $L-1$. Since
@@
×$\times\exp\left\{i\sum_{m^{\prime\prime}=1}^{M}\xi(\widetilde{q}^{(0)}_{x}(2m^{\prime\prime}-1))^{\star}e^{-i(2m^{\prime\prime}-1)\omega t}\right\}\mathrm{d}t.$ (5.3)
Denote by $C(1)=[z\in\mathbb{C}:\,|z|=1]$, a contour oriented counter-clockwise. Using contour integration we can rewrite the utmost right hand side of (5) as
@@
×$\times\prod_{m^{\prime\prime}=1}^{M}\sum_{n^{\prime}_{m^{\prime\prime}}=0}^{N}\frac{(i\xi(\widetilde{q}_{x}^{(0)}(2m^{\prime\prime}-1))^{\star})^{n^{\prime}_{m^{\prime\prime}}}}{(z^{2m^{\prime\prime}-1})^{n^{\prime}_{m^{\prime\prime}}}(n^{\prime}_{m^{\prime\prime}})!}\frac{\mathrm{d}z}{z^{2m+1}}$
@@
×$\times\delta\Big{(}n_{1}+3n_{2}+\ldots+(2M-1)n_{M}-2m-\big{(}n_{1}^{\prime}+3n_{2}^{\prime}+\ldots+(2M-1)n_{M}^{\prime}\big{)}\Big{)}$
×$\times\prod_{m^{\prime}=1}^{M}\frac{\big{(}\widetilde{q}_{x}^{(L-1)}(2m^{\prime}-1)\big{)}^{n_{m^{\prime}}}}{n_{m^{\prime}}!}\prod_{m^{\prime\prime}=1}^{M}\frac{\big{(}(\widetilde{q}_{x}^{(0)}(2m^{\prime\prime}-1))^{\star})^{n^{\prime}_{m^{\prime\prime}}}}{(n^{\prime}_{m^{\prime\prime}})!}.$
@@ 5.1. Remark about non-uniqueness
Assume that $V(\cdot)$ is of $C^{2}$ class of regularity. According to Remark 2.7, by adjusting $\omega$ to be sufficiently large, we can make $\overline{\nu}_{0}$ the range of the convergence of the perturbative scheme - as large as we wish. Then for some $|\nu|<\overline{\nu}_{0}$ the potential $W(q;\nu)$ could admit some other local minimum, say at $\overline{q}\not=0$. Then $W^{\prime}(\overline{q};\nu)=\omega_{0}^{2}\overline{q}+\nu V^{\prime}(\overline{q})=0$. Assume furthermore that $W^{\prime\prime}(\overline{q};\nu)=\omega_{0}^{2}+\nu V^{\prime\prime}(\overline{q})>0$. Then, obviously
@@
The only remaining part is the proof of strict positivity of the work functional. Suppose that ${\cal W}(\nu)=0$.
Then $p_{-N,{\rm p}}(t)=p_{N,{\rm p}}(t)\equiv 0$, $t\geqslant 0$. This implies that $q_{-N,{\rm p}}(t)\equiv q_{-N,{\rm p}}$, $q_{N,{\rm p}}(t)\equiv q_{N,{\rm p}}$, for all $t\geqslant 0$, and some constants $q_{-N,{\rm p}}$ and $q_{N,{\rm p}}$. Then $q_{-N+1,{\rm p}}(t)$ must solve the equation

### v4

@@ Abstract.
We study the propagation of energy in one-dimensional anharmonic chains subject to a periodic, localized forcing. For the purely harmonic case, forcing frequencies outside the linear spectrum produce exponentially localized responses, preventing equi-distribution of energy per degree of freedom. We extend this result to anharmonic perturbations with bounded second derivatives and boundary dissipation, proving that for small perturbations and non-resonant forcing, the dynamics converges to a periodic stationary state with energy exponentially localized uniformly in the system size. The perturbed periodic state is described by a convergent power type expansion in the strength of the anharmonicity. This excludes chaoticity induced by anharmonicity, independently of the size of the system. Our perturbative scheme can also be applied in higher dimensions.
Key words and phrases:
anharmonic chain, periodic force, spectrum of the infinite harmonic chain, perturbation series, stability.
2000 Mathematics Subject Classification:
80A19,80M22,82C22,82C70,70J35
@@ 1. Introduction
The approach to equilibrium which involves energy sharing among degrees of freedom of large systems with non-linear dynamics is one of the central problems in statistical mechanics. This motivated the numerical experiment in the seminal work of Fermi-Pasta-Ulam-Tsingou [18], which was perhaps the first application of digital computers to a non-military scientific goal. They considered a one-dimensional anharmonic chain of $N=32$ oscillators, initially prepared in a periodic low-energy configuration. Surprisingly, the simulated evolution failed to show thermalization, i.e. an equidistribution of the energy among all other frequencies generated by the nonlinearity of the dynamics. The results of these simulations remain a source of investigation and a challenge to theoreticians [7, 19, 16, 33]. Later studies repeated the simulations for much larger systems and at higher energies, showing that thermalization can indeed occur. However, no complete analytical results in this direction are known.
In this article we approach the problem from a different perspective. Instead of starting the dynamics of the anharmonic chain in a given periodic configuration, we apply a periodic force of period $\theta$ at a single site of the chain. For the linear dynamics it is known that only frequencies within the spectral interval can propagate through the system. In fact, in [20] we proved that if the forcing frequency lies outside the frequency spectrum of the system, the perturbation induced by the forcing remains exponentially localized. This is not surprising, since linear dynamics is completely integrable and cannot thermalize.
Here we consider anharmonic perturbations of the linear dynamics, with potentials having bounded second derivatives. Dissipation is applied at the boundaries, ensuring the existence of a periodic stationary state for any system size $N$. We prove that, for sufficiently small perturbations and if the forcing frequency is outside the resonance spectrum of the unperturbed linear system, the dynamics, starting from any initial configuration, converges in time to a periodic stationary state with energy exponentially localized uniformly in $N$. In particular, this implies that no thermalization occurs under these circumstances. This result is obtained via a rigorous perturbative scheme around the purely harmonic case. Although the method can be extended to higher dimension and infinite systems, we refrain from doing so here for simplicity.
An open question remains as to what happens when the forcing frequency lies within the resonance interval of the unperturbed linear system, or when the anharmonicity is unbounded (as in the FPUT case); see some numerical results in the companion article [21].
As far as we know, there are no other analytical results of this kind for non-linear dynamics that are uniform in the system size. There has also been considerable work on transport properties of anharmonic chains, both for systems coupled to thermal reservoirs and for systems driven by external periodic forces but at positive temperature, see [1, 2, 14, 5, 31, 11, 12, 13, 23, 24, 17, 35, 36] and references therein.
Computer simulations on similar systems, where a quartic anharmonicity is added to a pinned harmonic potential, reveal a rich variety of steady states emerging as the driving parameters of the external force are varied [27, 25, 5, 6, 34, 28, 8, 22].
2. Microscopic dynamics
@@
We denote ${\mathbb{Z}}_{N}:=\{-N,\ldots,N\}$. The microscopic dynamics of the chain are given by the forced Hamiltonian system with the Hamiltonian given by
${\cal H}_{N}({\bf q},{\bf p})=\sum_{x\in{\mathbb{Z}}_{N}}\left[\frac{p_{x}^{2}}{2}+\frac{1}{2}(q_{x}-q_{x-1})^{2}+\frac{\omega_{0}^{2}q_{x}^{2}}{2}+\nu\Big{(}V(q_{x})+U(q_{x}-q_{x-1})\Big{)}\right],$ (2.2)
friction on both endpoints of the lattice interval and a driving periodic force at $x=0$. This yields
@@
with strictly positive parameters $\gamma,\,\omega_{0},\,\theta$. The Neumann laplacian $\Delta$ and the discrete gradient $\nabla$ are defined as
@@
We prove that when all integer multiplicities of $\omega=\frac{2\pi}{\theta}$ are outside the interval containing the spectrum of the harmonic chain ${\cal I}:=[\omega_{0},\omega_{u}]$, with $\omega_{u}:=\sqrt{\omega_{0}^{2}+4}$, a unique state of the system is achieved, as $t\to+\infty$, for an arbitrary initial condition. In addition, the state is $\theta$-periodic and given by a perturbative expansion in the parameter $\nu$, that converges for $|\nu|<\nu_{0}$, uniformly in the size of the system $N$. More precisely
@@
where $\delta_{*}:=\inf\Big{[}|(m\omega)^{2}-w^{2}|:\,m\in{\mathbb{Z}},\,w\in{\cal I}\Big{]}$, and ${\mathfrak{V}}=\|V^{\prime\prime}\|_{\infty}+3\|U^{\prime\prime}\|_{\infty}$.
Finding a periodic stable solution of (2.3) by a convergent perturbative series is, we believe, unlikely for nonlinear anharmonicity that grows faster than quadratic, considered in [3, 30, 16], and, as far as we know, not proven for any other non-equilibrium system. Our results do not extend to the case where some integer multiplicity of $\omega$ lies inside spectrum of the harmonic chain or $|\nu|>\nu_{0}$.
@@ 3.4. Global stability of the unique periodic solution in the anharmonic case
Theorems 3.2 and 3.3 hold even in the absence of dissipation at the boundaries ($\gamma=0$) in the case of an infinite chain ($N=+\infty$). Of course in the case $\gamma=0$ the unique periodic solution has no stability properties (as can be seen even for $\nu=0$), it is just a very special solution of (2.3). Stability holds only when $\gamma>0$.
@@ 3.5. Work done by periodic forcing
${{\cal W}_{N}}(\nu)=\frac{1}{\theta}\int_{0}^{\theta}{\cal F}(t/\theta)p_{0,{\rm p}}(t;\nu)\mathrm{d}t.$ (3.18)
@@
Suppose that $\gamma>0$ and $\big{(}q_{x,{\rm p}}(\cdot;\nu)\big{)}$ is a $\theta$-periodic solution of (2.3). Furthermore assume that the potential $U$ is such that the equation222 Condition 3.20 means that the interaction potential $\frac{1}{2}r^{2}+\nu U(r)$, between the neighboring atoms does not become ”flat” in some intervals, which would prevent the energy transfer in the chain.
@@
${\cal W}_{N}(\nu)=\frac{\gamma}{\theta}\Big{(}\int_{0}^{\theta}p_{-N,{\rm p}}^{2}(s)\mathrm{d}s+\int_{0}^{\theta}p_{N,{\rm p}}^{2}(s)\mathrm{d}s\Big{)}>0.$ (3.21)
In addition,
$\lim_{N\to+\infty}{\cal W}_{N}(\nu)=0.$ (3.22)
The proof of this result is presented in Section 7.
@@ 3.6. Spatial decay of the periodic solution as $N\to\infty$
There exist positive constants $C_{1},A,\rho$ depending only on $\omega,\gamma$ and independent of $N$ such that for $|\nu|<\nu_{0}\wedge C_{1}$
@@ 3.7. The case of odd harmonic modes: loss of uniqueness
We stress here the fact that Theorem 3.7 does not claim either the uniqueness, or stability of the solution given by the series (3.15), in the case $\nu_{0}\leqslant|\nu|<\overline{\nu}_{0}$. In this interval the periodic solution defined by (3.15) still exists but it may be neither stable nor unique. As we shall see in Section 6.1, the system (2.3) may admit more than one $\theta$-periodic solution. Obviously, in that case, the stability in the sense claimed in Theorem 3.4 cannot hold.
@@ 4. The case $N=0$: a single anharmonic oscillator with damping
[…]
@@ 5.1. Time harmonics of a periodic solution
$0$ $=\big{[}(\omega m)^{2}-\omega_{0}^{2}+\Delta-i\gamma\omega m(\delta_{x,-N}+\delta_{x,N})\big{]}\widetilde{q}_{x}(m;\nu)+\widehat{\rm F}_{m}\delta_{x,0}$ (5.4)
@@ 5.2. The case of a harmonic chain
$\begin{split}\widetilde{q}_{x}^{(L)}(m;\nu)=\frac{1}{\theta}\int_{0}^{\theta}e^{-im\omega t}q_{x,{\rm p}}^{(L)}(t;\nu)dt,\quad m\in\mathbb{Z}_{N}.\end{split}$ (5.12)
@@
$0$ $=\left[\Delta+(m\omega)^{2}-\omega_{0}^{2}-i\gamma m\omega(\delta_{-N}(x)+\delta_{N}(x))\right]\widetilde{q}_{x}^{(L)}(m)-\;\widetilde{v}_{x,L-1}(m;\nu),\;\quad x\in{\mathbb{Z}}_{N},$ (5.13)
@@ 5.4. Regularity of the periodic solution
$0$ $=\Big{(}\Delta+(m\omega)^{2}-\omega_{0}^{2}-i\gamma m\omega(\delta_{-N}+\delta_{N})\Big{)}\Big{(}\delta\widetilde{Tf}_{x}(m)\Big{)}+$ (5.33)
@@
By the Banach contraction principle there exists a unique solution of the equation Tf=f{Tf}=f in $L^{2}\big{(}[0,\theta];{\mathbb{R}}^{2N+1}\big{)}$. Then $f(t)$ given by (5.32) is the unique $\theta$-periodic solution of (2.3). Hence the conclusion of the theorem follows. ∎
@@
see (5.14). To see (6.2) (and (6.1)) assume for simplicity that $U\equiv 0$. The argument in the general case is similar. Since $V^{\prime}(\cdot)$ is odd and real valued, its Fourier transform $\widehat{V^{\prime}}(\xi)$ is also an odd function. To further simplify the argument we suppose that its support is compact. The general case can be treated by an approximation argument. Thanks to (5.9) and the fact that $\widehat{F}_{2m}=0$ for all $m$, we conclude that (6.1) holds for $L=0$. Suppose now that (6.2) is true for some non-negative integer L1L-1. Since
@@
$\times\exp\left\{i\sum_{m^{\prime\prime}=1}^{M}\xi{(\widetilde{q}^{(L-1)}_{x}(2m^{\prime\prime}-1))^{\star}}e^{-i(2m^{\prime\prime}-1)\omega t}\right\}\mathrm{d}t.$ (6.3)
margin: Corr.!
Denote by $C(1)=[z\in\mathbb{C}:\,|z|=1]$, a contour oriented counter-clockwise. Using contour integration we can rewrite the utmost right hand side of (6) as margin: Corr.!
@@
$\times\prod_{m^{\prime\prime}=1}^{M}\sum_{n^{\prime}_{m^{\prime\prime}}=0}^{N}\frac{(i\xi{(\widetilde{q}_{x}^{(L-1)}(2m^{\prime\prime}-1))^{\star}})^{n^{\prime}_{m^{\prime\prime}}}}{(z^{2m^{\prime\prime}-1})^{n^{\prime}_{m^{\prime\prime}}}(n^{\prime}_{m^{\prime\prime}})!}\frac{\mathrm{d}z}{z^{2m+1}}$
@@
$\times\delta\Big{(}n_{1}+3n_{2}+\ldots+(2M-1)n_{M}-2m-\big{(}n_{1}^{\prime}+3n_{2}^{\prime}+\ldots+(2M-1)n_{M}^{\prime}\big{)}\Big{)}$
$\times\prod_{m^{\prime}=1}^{M}\frac{\big{(}\widetilde{q}_{x}^{(L-1)}(2m^{\prime}-1)\big{)}^{n_{m^{\prime}}}}{n_{m^{\prime}}!}\prod_{m^{\prime\prime}=1}^{M}\frac{\big{(}{(\widetilde{q}_{x}^{(L-1)}(2m^{\prime\prime}-1))^{\star}})^{n^{\prime}_{m^{\prime\prime}}}}{(n^{\prime}_{m^{\prime\prime}})!}.$
@@ 6.1. Remark about non-uniqueness
Assume that $V(\cdot)$ is of $C^{2}$ class of regularity. According to Remark 3.7, by adjusting $\omega$ to be sufficiently large, we can make $\overline{\nu}_{0}$ - the range of the convergence of the perturbative scheme - as large as we wish. Then for some $|\nu|<\overline{\nu}_{0}$ the potential $W(q;\nu)$ could admit some other local minimum, say at $\overline{q}\not=0$. Then $W^{\prime}(\overline{q};\nu)=\omega_{0}^{2}\overline{q}+\nu V^{\prime}(\overline{q})=0$. Assume furthermore that $W^{\prime\prime}(\overline{q};\nu)=\omega_{0}^{2}+\nu V^{\prime\prime}(\overline{q})>0$. Then, obviously
@@
7.1. Proof of positivity of the work functional
The fact that ${\cal W}_{N}(\nu)\geqslant 0$ is a consequence of the identity (3.19). The only remaining part is the proof of strict positivity of the work functional. Suppose that ${\cal W}_{N}(\nu)=0$. Then $p_{-N,{\rm p}}(t)=p_{N,{\rm p}}(t)\equiv 0$, $t\geqslant 0$. This implies that $q_{-N,{\rm p}}(t)\equiv q_{-N,{\rm p}}$, $q_{N,{\rm p}}(t)\equiv q_{N,{\rm p}}$, for all $t\geqslant 0$, and some constants $q_{-N,{\rm p}}$ and $q_{N,{\rm p}}$. Then $q_{-N+1,{\rm p}}(t)$ must solve the equation